# PRISM: Leveraging Prototype Patient Representations with Feature-Missing-Aware Calibration for EHR Data Sparsity Mitigation

### v4

@@ 6.3 Observations and Analysis
Table 7 showcase PRISM’s data preparation time in preprocessing and the overall time for training&validation, and testing on MIMIC datasets. The PRISM pipeline seamlessly computes feature missing statuses in the LOCF pipeline with little extra cost. Moreover, the overall pipeline can be completed in 10 minutes.
@@
Cross-Feature Attention Map
@@
Figure 4: Cross-Feature Attention. Without (Left) / With (Right).
Feature Decay Rates Observation
@@
Figure 5: Feature decay rates observation
@@ 7 Conclusions
This study, involving the analysis of EHR data is committed to upholding high ethical standards. The employed datasets are all de-identified, ensuring patient confidentiality and privacy.

### v5

@@ 6.3.4 Model Efficiency and Complexity
Table 7 presents the data preparation time for PRISM on MIMIC-III and MIMIC-IV datasets, including preprocessing, training, validation, and testing. The PRISM pipeline seamlessly computes feature missing statuses in the LOCF pipeline with little extra cost. Moreover, the entire pipeline can be completed within 10 minutes for each dataset.
@@
Dataset Challenge-2012 eICU
Prep. w/o Impute LOCF +Ours w/o Impute LOCF +Ours
Time 18.46 56.59 85.14 52.99 146.54 170.21
Pipeline Data Prep. Train+Val. Test Data Prep. Train+Val. Test
Time 85.14 85.55 1.47 170.21 268.17 4.99
6.3.5 Cross-Feature Attention Map
@@
Figure 4: Cross-feature attention maps from PRISM: Without (Left) / With (Right) feature confidence calibration. The maps use data from a single MIMIC-IV patient record to show PRISM’s reduction in attention to unreliable features (capillary refill rate, fraction inspired oxygen, height) due to high missing rates.
6.3.6 Patient Representation Visualization
To investigate the impact of the missing-feature-calibration process on the hidden representations of patient data, we apply t-SNE to project these representations onto a two-dimensional space using the test set of all patients from the MIMIC-IV dataset. Figure 5 illustrates the t-SNE embeddings of the patient representations generated by PRISM, both with and without the application of the missing-feature-calibration process. The calibrated representations (Figure 5(b)) exhibit improved separation and compactness, particularly for patients with mortal outcomes, compared to the representations without calibration (Figure 5(a)). This observation suggests that the missing-feature-calibration process enables PRISM to learn more informative and discriminative representations of patient data by effectively handling missing features and capturing the underlying patterns and relationships within the data.
(a) Without calibration
(b) With calibration
Figure 5: t-SNE visualization of patient representations from PRISM. (a) shows the embeddings without the missing-feature-calibration process, and (b) depicts embeddings with the process. (b)’s representations are more compact among dead outcome patients, showcasing it learns better representations.
6.3.7 Feature Decay Rates Observation
@@
Figure 6: Adaptive learning feature decay rates. The graph shows varying decay rates: high for acute-indicator features like heart rate and pH, and low for longer-term relevant features like sex and systolic blood pressure.
7 Limitations and Further Work
We propose potential limitations of our current model and suggest directions for future research:
  • Fairness Concerns: Like many machine learning applications, it is essential to evaluate our model’s fairness across various demographic groups [7]. Discovering bias issue in similar patients cohort could be further explored.
  • Scalability Issues: The proposed model’s scalability to more massive datasets or integration within real-time healthcare systems is yet to be assessed.
  • Prototype Patient Representations: Though leveraging prototype patient representations has shown promise, further work is needed to understand the diversity of these prototypes. More intricate mechanisms for prototype generation, beyond similarity metrics, could also be explored.
@@ 8 Conclusions
This study, which involves analyzing Electronic Health Records (EHR) from the MIMIC-III, MIMIC-IV, PhysioNet Challenge 2012, and eICU datasets, is dedicated to maintaining high ethical standards. All datasets used are de-identified to ensure patient confidentiality and privacy. By signing the data use agreement, we reaffirm our commitment to responsible data handling and usage. Our methodology is designed to minimize potential harm and to produce findings that are as unbiased and equitable as possible, reflecting the diverse and complex nature of medical data. We have also taken rigorous steps to ensure that our research adheres to these ethical values.